OpenStudy (anonymous):

The pH of a substance is given by the formula pH=-log[H^+], where [H^+] is the concentration of hydrogen ions and log x means log(10)x. Find the approximate concentration of hydrogen ions in a substance that has a pH of 6.8. ****use scientific notation

OpenStudy (anonymous):

oh i remember messing this one up before

OpenStudy (anonymous):

lol, yeah, as have I

OpenStudy (anonymous):

\[6.8=-\log(H^+)\] \[-6.8=\log(H^+)\] \[10^{-6.8}=H^+\]

OpenStudy (anonymous):

now we have to write it in scientific notation!

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=10^-6.8

OpenStudy (anonymous):

even gives it to you in scientific notation. it is \[1.585\times 10^{-7}\] rounded